A total of 550 tickets were sold adults tickets cost 2.00 children tickets cost 1.50 a total sold for 950.00 how many were adult and how many were children?
Answer by ptaylor(2198)

Let x=number of children's tickets sold
then 550-x=number of adult tickets sold
1.50*x+2.00(550-x)=950
1.5x+1100-2x=950 subtract 1100 from each side
-0.5x=-150
x=300----------------------number of children's tickets sold
550-x=550-300=250-----number of adult tickets sold
CK
300(1.50)+250(2.00)=950
450+500=950
950=950
another way:)
x=children's
y=adults
x+y=550
1.50x+2.00y=950
multiply first equation by 2, then subtract 2nd from 1st
0.5x=150
x=300
etc.
Hope this helps----ptaylor